Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 11

1. Section 5.1, problem 22:

Figure 1: Section 5.1, problem 22(a,b)

Figure 2: Section 5.1, problem 22(c)

2. Section 5.3, problem 87 (optional):

Figure 3: Section 5.3, problem 87 3